If the length of a rectangle is 8 inches and the width is 1/2 the size of the length , what are the rectangles area and perimeter?

Answer by nic0le116(25)

Area= Length * width
You are given the length as 8.
Since it states that the width is half of the length, this means the width is 4.
Area= 8*4=32in
-------------------------------------------
Perimeter= the sum of all the sides.
P=8+8+4+4=24 inches
---------------------------------------------
ANSWERS=
Area= 32in
Perimeter=24in